PerfectlySoft Corp. is experiencing rapid growth. Dividends are expected to grow at 29 percent per year during the next three years, 13 percent over the following year, and then 2 percent per year thereafter indefinitely. The required return on this stock is 9.42 percent, and the stock currently sells for $71.58 per share. What is the projected dividend (in $) for the coming year? Answer to two decimals, carry intermediate calcs. to four decimals.
Answer #1

P0 = [{D0 * (1 + g1)} / (1 + r)] + [{D0 * (1 + g1)2} / (1 + r)2] + [{D0 * (1 + g1)3} / (1 + r)3] + [{D0 * (1 + g1)3 * (1 + g2)} / (1 + r)4] + [{D0 * (1 + g1)3 * (1 + g2) * (1 + gC)} / {(r - gC) * (1 + r)4}]

$71.58 = [{D0 * (1 + 0.29)} / (1 + 0.0942)] + [{D0 * (1 + 0.29)2} / (1 + 0.0942)2] + [{D0 * (1 + 0.29)3} / (1 + 0.0942)3] + [{D0 * (1 + 0.29)3 * (1 + 0.13)} / (1 + 0.0942)4] + [{D0 * (1 + 0.29)3 * (1 + 0.13) * (1 + 0.02)} / {(0.0942 - 0.02) * (1 + 0.0942)4}]

$71.58 = [1.1789 * D0] + [1.3899 * D0] + [1.6386 * D0] + [1.6922 * D0] + [23.2625 * D0]

$71.58 = [29.1621 * D0]

D0 = $71.58 / 29.1621 = $2.4546

D1 = D0 * (1 + g1) = $2.4546 * (1 + 0.29) = $3.17